7x + 2y = 0
4y = -14x
what are the steps to solve
is this infinitely many solutions or no solution?

Answer by jorel1380(3719)

7x+2y=0
4y=-14x
14x+4y=0
14x-14x=0
There are infinitely many solutions to this..